AN ESTIMATE FOR THE GAUSS CURVATURE OF MINIMAL

SURFACES IN ℝ^{⋗} WHOSE GAUSS MAP OMITS A SET OF HYPERPLANES

ROBERT OSSERMAN MIN RU

Abstract. We give an estimate of the Gauss curvature for minimal surfaces in ℝ^{m} whose Gauss map omits more than m(m + 1)/2 hyperplanes in ℙ^{m−1}(ℂ).